OpenStudy (anonymous):

Find all the roots of \[z^{3}=\sqrt{3}+i\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[z^3 = \sqrt3 + i \] \[\theta=\tan^{-1}\frac{y}{x} \implies \frac{1}{\sqrt3}\]

OpenStudy (anonymous):

\[\left\{z\to \left(i+\sqrt{3}\right)^{1/3},z\to -(-1)^{1/3} \left(i+\sqrt{3}\right)^{1/3},z\to (-1)^{2/3} \left(i+\sqrt{3}\right)^{1/3}\right\} \]